# DiffX: Guide Your Layout to Cross-Modal Generative Modeling

### v1

@@ Abstract
Diffusion models have made significant strides in text-driven and layout-driven image generation. However, most diffusion models are limited to visible RGB image generation. In fact, human perception of the world is enriched by diverse viewpoints, including chromatic contrast, thermal illumination, and depth information. In this paper, we introduce a novel diffusion model for general layout-guided cross-modal “RGB+X” generation, called DiffX. We firstly construct the cross-modal image datasets with text descriptions using the LLaVA model for image captioning, supplemented by manual corrections. Notably, DiffX presents a simple yet effective cross-modal generative modeling pipeline, which conducts diffusion and denoising processes in the modality-shared latent space, facilitated by our Dual-Path Variational AutoEncoder (DP-VAE). Furthermore, we incorporate the gated cross-attention mechanism to connect the layout and text conditions, leveraging Long-CLIP for embedding long captions to enhance user guidance. Through extensive experiments, DiffX demonstrates robustness and flexibility in cross-modal generation across three RGB+X datasets: FLIR, MFNet, and COME15K, guided by various layout types. It also shows the potential for adaptive generation of “RGB+X+Y” or more diverse modalities. Our code and processed image captions are available at https://github.com/zeyuwang-zju/DiffX.
@@ 1 Introduction
Over the past decade, there have been significant advancements in image generation, driven by the improvement of deep generative models such as Variational AutoEncoders (VAEs) (Kingma and Welling 2013) and Generative Adversarial Networks (GANs) (Goodfellow et al. 2014). Notably, the recent diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) like DALL-E (Ramesh et al. 2021, 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (SD) (Rombach et al. 2022) have gained popularity for their ability to generate high-fidelity images conditioned on text descriptions or class labels. Moreover, the field has witnessed the rise of layout-to-image models based on various layout conditions, including bounding boxes, semantic maps, and keypoints (Xie et al. 2023; Li et al. 2023; Ma et al. 2024). The success of diffusion models can be attributed to its stable learning objective and the availability of large-scale multi-modal training data.
@@
Figure 2: (a) The process of preparing the image captions. (b) Examples of cross-modal images, labels, and prepared captions.
Current layout-to-image diffusion models primarily focus on generating high-quality visible RGB images. However, our perception of the world is enriched by various perspectives beyond the visible spectrum, such as thermal illumination and depth information. Cross-modal visual understanding typically leverages the input format of “RGB+X”, where X represents additional data like Thermal (T) or Depth (D) images. It offers a more comprehensive representation compared with RGB-only understanding, especially in challenging environments. At present, a significant limitation lies in the scarcity of training data containing pixel-aligned RGB+X image pairs. If the conventional generative models are adopted for cross-modal data augmentation, we can only generate RGB and X images separately, and as a result, the RGB and X images cannot form consistent pairs of matching data. Consequently, we wonder if we can generate the cross-modal images simultaneously under layout guidance by an integrated model?
In this work, we introduce a novel diffusion model for cross-modal image generation, called DiffX. As shown in Figure 1, the proposed DiffX is capable of generating RGB+X image pairs guided by various layouts with user instructions. Our main contributions are as follows:
  • Due to the limited cross-modal images with text descriptions, we leverage the Large-Multimodal Model (LMM) to prepare the captions, followed by manual corrections.
  • We propose an effective cross-modal generative modeling pipeline, which performs the diffusion and denoising processes in the modality-shared latent space. Specially, the DP-VAE is implemented for cross-modal encoding.
  • We employ the gated cross-attention mechanism to establish a comprehensive connection between layout and text conditions, where the Long-CLIP (Zhang et al. 2024) is adopted to capture critical information in captions.
  • Experiments demonstrate that DiffX is capable of generating high-quality and coherent RGB+X image pairs based on bounding boxes, semantic maps, salient maps, and edge contours. Additionally, it demonstrates adaptability to “RGB+X+Y” or diverse-modality generation.
@@ 2.1 Cross-Modal Visual Understanding
Cross-modal visual understanding normally adopts RGB+X input data. Thermal images have shown advantages in detecting objects in low-light conditions. For instance, CFT (Fang, Han, and Wang 2021) and ICAFusion (Shen et al. 2024) incorporate Transformer-based RGB+T feature fusion modules for multispectral object detection. In semantic segmentation, EGFNet (Zhou et al. 2022) leverages prior edge maps of RGB+T images to capture detailed contour information. Secondly, depth images offer valuable insights into the 3D layouts and spatial structures. For example, HidaNet (Wu et al. 2023) utilizes a granularity-based attention mechanism to enhance the advantages of RGB+D features for Salient Object Detection (SOD). However, current cross-modal visual understanding is hindered by the lack of well-aligned RGB+X image pairs for training deep models.
@@ 2.2 Layout-to-Image Generation
Based on Denoising Diffusion Probabilistic Model (DDPM) (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020), diffusion models have undergone advancements in training and sampling techniques (Song, Meng, and Ermon 2021; Nichol and Dhariwal 2021). Recent studies have proposed layout-to-image diffusion models, which provide precise instructions for the object placement. Notably, LAW-Diffusion (Yang et al. 2023) integrates a spatial parser and adaptive guidance, enabling complex scene generation. BoxDiff (Xie et al. 2023) directly incorporates spatial conditions into the training-free denoising process. InstanceDiffusion (Wang et al. 2024) focuses on precise instance-level control in image generation.
@@ 2.3 Multi-Modal Generation
Multi-modal generation has become a challenging task and attracted widespread research. MM-Diffusion (Ruan et al. 2023) integrates featuring two interconnected denoising autoencoders and a sequential multi-modal U-Net architecture for joint audio-video generation. Moreover, MM-Interleaved (Tian et al. 2024) dynamically extracts information from multi-scale features for image-text generation. In addition, ControlVAR (Li et al. 2024), MT-Diffusion (Chen et al. 2024), and 4M (Mizrahi et al. 2024) adopt large-scale natural datasets or private datasets for end-to-end training on conditional generation. Despite these advancements, there remains a notable gap in generative models focused on layout-guided cross-modal “RGB+X” image generation.
3 LMM-Assisted Caption Generation
For training our diffusion model for cross-modal generation, it is necessary to prepare detailed captions for images, where the workflow is illustrated in Figure 2 (a).
In detail, the first step involves the utilization of the advanced LMM, namely LLaVA-v1.5-7b model (Liu et al. 2024), to extract descriptive information related to weather conditions, environmental settings, surrounding elements, transportation conditions, and identifiable objects within the RGB images. Despite the effectiveness of LLaVA in capturing visual content, it is prone to inaccuracies and omissions captioning, especially in complex scenes. Additionally, it lacks the ability to simultaneously annotate complementary information in both RGB and X images. Therefore, a correction procedure is implemented after the generation of raw captions. In cases where the generated captions exhibit minimal errors, manual corrections are applied to enhance accuracy. However, instances of significant errors or insufficient descriptions prompting ambiguity necessitate further corrections. Under such conditions, the GPT-4 language model (Achiam et al. 2023) serves as a valuable tool for providing supplementary descriptions and rectifying inaccurate information to ensure a comprehensive image depiction. By combining the capability of the LLaVA model with the manual correction and assistance of GPT-4, the captioning process achieves a high level of fidelity and accuracy in conveying the visual context in the cross-modal datasets.
@@ 4 DiffX for Cross-Modal Generation
In this section, we introduce our proposed DiffX, designed to generate cross-modal “RGB+X” images, as illustrated in Figure 3. We first introduce the preliminaries on Latent Diffusion Model (LDM). Then, we describe the workflow and components of our DiffX in detail. Finally, we provide a discussion on the advantages and unique features of DiffX.
@@ 4.1 Preliminaries on Latent Diffusion Model
Building upon DDPM, LDM and the successor SD (Rombach et al. 2022) have pushed the boundaries of diffusion models by using a low-dimensional latent space. Inside LDM, a prior VAE is employed to extract the latent representation $\boldsymbol{z}$ of the ground-truth image $\boldsymbol{I}$. Subsequently, the diffusion model $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$, normally implemented as a UNet (Ronneberger, Fischer, and Brox 2015) with residual (He et al. 2016) and self-attention (Vaswani et al. 2017) blocks, is adopted to denoise the anormally-distributed $\boldsymbol{\epsilon}$ in the noisy latent $\boldsymbol{z}_{t}$ at time step $t$. The training objective can be represented as:
@@
where $\boldsymbol{c}$ denotes the conditional input, typically a text embedding generated by a text encoder like CLIP (Radford et al. 2021) in text-to-image generation.
Following the training process, the diffusion model gradually generates $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ from a randomly initialized noisy latent $\boldsymbol{z}_{T}$. Finally, the decoder of VAE is employed to generate the desired image $\boldsymbol{I}^{\prime}$ based on $\boldsymbol{z}_{0}$.
@@ 4.2 Cross-Modal Modeling Formulation
In our DiffX model, we firstly obtain the joint distribution by mapping the “RGB+X” modalities into the shared latent space. Specifically, during the forward process, the noising transition from $t-1$ to $t$ is conditioned on the joint modality $\mathbf{M}=\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$, which can be expressed as:
@@
Meanwhile, the reverse process can be regarded as the reverse of the forward distributions, resulting in a joint distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1},\mathbf{M}|\boldsymbol{z}_{t})$ at time step $t$:
@@
where $\boldsymbol{\theta}$ denotes the parameters of the our DiffX model.
@@ 4.3 Dual-Path Variational AutoEncoder
As illustrated in Figure 4, the proposed Dual-Path Variational AutoEncoder (DP-VAE) is critical for the cross-modal generation task. In contrast to the conventional VAE used in LDM, our DP-VAE employs the form of “RGB+X” input and “RGB+X” output. In detail, the DP-VAE utilizes a single encoder $\mathcal{E}$ to encode the input $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$} (denoted as $\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$ in Eq. (2)) into a latent representation $z$, and incorporates two parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ to generate the corresponding output $\boldsymbol{I}_{v}^{\prime},\boldsymbol{I}_{x}^{\prime}$}. The input images are processed through the separate convolutional layers, followed by element-wise addition before being fed to $\mathcal{E}$.
In addition, we have observed that image modality “X” usually exhibits distinct contour edges and prominent object positions. To enhance the reconstruction ability of the DP-VAE framework, we apply a Laplacian Pyramid (LP) (Burt and Adelson 1987) to extract the frequency-aware information from the input cross-modal image pairs. The extracted features are embedded into the various scale layers of the encoder through convolutional layers, enabling a comprehensive enhancement across various scales.
4.4 DiffX Model in Modality-Shared Space
The overall workflow of the DiffX model is shown in Figure 3. The noisy latent $\boldsymbol{z}_{T}$ is randomly initialized, which is adopted for generating $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ by our DiffX-UNet. The diffusion processes involve grounding the effective conditional input, including the text captions and layout conditions. Finally, the parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ from the pre-trained DP-VAE are employed to generate the cross-modal “RGB+X” images based on $\boldsymbol{z}_{0}$, which is represented as the denoised joint distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{0},\mathbf{M}|\boldsymbol{z}_{1})$ in Eq. (3).
@@ 4.4 DiffX Model in Modality-Shared Space
Before the diffusion process, it is necessary to embed the layout conditions into grounding features to be fed into the DiffX-UNet.
1) For the box-based layout, we embed each box $\boldsymbol{b}_{i}=[x_{\text{min}},y_{\text{min}},x_{\text{max}},y_{\text{max}}]$ with label $\boldsymbol{l}_{i}$ into one grounding token. In detail, $\boldsymbol{b}_{i}$ and $\boldsymbol{l}_{i}$ are embedded via Fourier mapping $\mathcal{F}(\cdot)$ (Tancik et al. 2020) and the pre-trained CLIP text encoder $f_{\text{clip}}(\cdot)$, respectively. Afterwards, a Multi-Layer Perception (MLP) with parameters $\phi$ is adopted to encode the concatenated features:
@@
2) For the Seg., SOD, and edge layout $\boldsymbol{m}$, the pre-trained ConvNeXtψ model (Liu et al. 2022) is employed to extract the map features. After that, the Position Embedding $\boldsymbol{p}$ is added. Finally, an MLP with parameters $\varphi$ is used to generate the layout conditional feature:
@@
where $\mathcal{R}$ denotes the reshaping process to transform the feature into flattened tokens.
@@
In our cross-modal generation task, the layout-based condition plays a crucial role in providing the general outline of objects within the images. However, text caption conditions $\boldsymbol{c}$ remain equally significant in cross-modal generation as they offer detailed descriptions of the images, encompassing aspects such as weather conditions, environments, surroundings, transportation modes, and more. Specifically, our task necessitates long text descriptions to capture the multitude of small objects and complexities present in the scene, as exemplified in Figure 2 (b). In our DiffX, we utilize the recent Long-CLIP model (Zhang et al. 2024) as the text encoder $f_{\text{L-clip}}(\cdot)$ to embed the text captions, which represents a significant advancement over its predecessor, CLIP, due to its ability to support long-text input. In detail, each caption $\boldsymbol{c}$ is embedded into a sequence of text embeddings $\boldsymbol{c}^{*}=f_{\text{L-clip}}(\boldsymbol{c})$.
@@
DiffX-UNet.
DiffX-UNet serves as the denoising model, featuring a combination of cascaded residual and Spatial-Transformer blocks. To enhance performance and efficiency, pre-trained weights from SD are employed for the frozen layers. The Spatial-Transformer block is illustrated in Figure 3, which is the fundamental component of the DiffX-UNet. It is composed of cascaded self-attention, cross-attention, and Feed Forward (FF) layers. The Feed Forward (FF) layer is an MLP with the middle dimensional expansion and skip connection. For the trainable FF layers in DiffX model, we add the scaling factors with the tanh activation. The cross-attention mechanism, derived from the self-attention mechanism, is formulated as follows:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}\right)\cdot V,$ (6)
where $Q$, $K$, and $V$ are derived from linear transformations of the features, and $d_{k}$ denotes the dimension of $K$. In our model, the visual tokens always serve as the modality to generate $Q$ in cross-attention layers.
In specific, the gated cross-attention layer and one FF layer are implemented prior to the Spatial-Transformer blocks, which aim to generate the enhanced layout conditional feature $\boldsymbol{H}^{*}$:
$\boldsymbol{H}^{\prime}=\boldsymbol{H}+\lambda\cdot\text{tanh}(\gamma_{1})\cdot\text{Cross-Attn}(\boldsymbol{H},\boldsymbol{c}^{*}),$ (7)
where the hyperparameter $\lambda$ plays a crucial role in balancing quality and controllability. Additionally, $\gamma_{1}$ and $\gamma_{2}$ are two adaptive learnable scalars inside tanh activation.
Afterwards, the resulting $\boldsymbol{H}^{*}$ is fed into the gated self-attention within the Spatial-Transformer blocks. The gated self-attention, inspired by GLIGEN (Li et al. 2023), acts as an adapter module to effectively capture cross-modal relationships. Similarly, an FF is layer is integrated after each gated self-attention to build the long-range interaction:
$\boldsymbol{T}^{\prime}=\boldsymbol{T}+\mu\cdot\text{tanh}(\delta_{1})\cdot\mathcal{T}(\text{Self-Attn}([\boldsymbol{T},f_{\text{linear}}(\boldsymbol{H}^{*})])),$ (9)

### v2

@@ Abstract
Diffusion models have made significant strides in language-driven and layout-driven image generation. However, most diffusion models are limited to visible RGB image generation. In fact, human perception of the world is enriched by diverse viewpoints, including chromatic contrast, thermal illumination, and depth information. In this paper, we introduce a novel diffusion model for general layout-guided cross-modal “RGB+X” generation, called DiffX. Firstly, we construct the cross-modal image datasets with text description by using LLaVA for image captioning, supplemented by manual corrections. Notably, DiffX presents a simple yet effective cross-modal generative modeling pipeline, which conducts diffusion and denoising processes in the modality-shared latent space, facilitated by our Dual-Path Variational AutoEncoder (DP-VAE). Moreover, we introduce the joint-modality embedder, which incorporates a gated cross-attention mechanism to link layout and text conditions. Meanwhile, the advanced Long-CLIP is employed for long caption embedding to improve user guidance. Through extensive experiments, DiffX demonstrates robustness and flexibility in cross-modal generation across three RGB+X datasets: FLIR, MFNet, and COME15K, guided by various layout types. It also shows the potential for adaptive generation of “RGB+X+Y” or more diverse modalities. Our code and constructed cross-modal image datasets are available at https://github.com/zeyuwang-zju/DiffX.
@@ 1 Introduction
Over the past decade, there have been significant advancements in image generation, driven by the improvement of deep generative models such as Variational AutoEncoders (VAEs) (Kingma and Welling 2013) and Generative Adversarial Networks (GANs) (Goodfellow et al. 2014). Notably, the recent diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) like DALL-E (Ramesh et al. 2021, 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (SD) (Rombach et al. 2022) have gained popularity for their ability to generate high-fidelity images conditioned on text description or class labels. Moreover, the field has witnessed the rise of layout-to-image models based on various layout conditions, including bounding boxes, semantic maps, and keypoints (Xie et al. 2023; Li et al. 2023; Ma et al. 2024). The success of diffusion models can be attributed to their stable learning objective and the availability of large-scale multi-modal training data.
@@
Figure 2: (a) The process of constructing the image captions. (b) Examples of cross-modal images, labels, and prepared captions.
Current layout-to-image diffusion models primarily focus on generating high-quality visible RGB images. However, our perception of the world is enhanced by a multitude of perspectives that extend beyond the visible spectrum, including thermal imaging and depth perception. Cross-modal visual understanding typically leverages the input format of “RGB+X”, where X represents additional data like Thermal (T) or Depth (D) images. It offers a more comprehensive representation compared with RGB-only understanding, especially in challenging environments. At present, a significant limitation lies in the scarcity of cross-modal data containing pixel-aligned RGB+X image pairs. If the conventional generative models are adopted for cross-modal data augmentation, we can only generate RGB and X images separately, and as a result, the RGB and X images cannot form consistent pairs of matching data. Consequently, we wonder if we can generate the cross-modal images simultaneously under layout guidance by an integrated model?
In this work, we introduce a novel diffusion model for cross-modal image generation, called DiffX. As shown in Figure 1, our proposed DiffX model is capable of generating RGB+X image pairs guided by various layouts with user instructions. Our main contributions are as follows:
  • We propose an effective cross-modal generative modeling pipeline, which performs the diffusion and denoising processes in modality-shared latent space. In particular, the DP-VAE is implemented for cross-modal encoding.
  • We propose a joint-modality embedder to establish the comprehensive connection between layout and text conditions via the gated cross-attention mechanism. Besides, we employ the advanced Long-CLIP (Zhang et al. 2024) to capture critical information in cross-modal captions.
  • Due to the limited cross-modal data with text descriptions, we leverage a Large-Multimodal Model (LMM) to construct captions, followed by manual corrections.
  • Our experiments demonstrate that DiffX can generate high-quality and coherent “RGB+X” image pairs based on bounding boxes, semantic maps, salient maps, and edge contours. Additionally, it demonstrates adaptability to “RGB+X+Y” or diverse-modality generation.
@@ 2.1 Cross-Modal Visual Understanding
Cross-modal visual understanding normally adopts RGB+X input data. Thermal images have shown advantages in detecting objects in low-light conditions. For instance, CFT (Fang, Han, and Wang 2021) and ICAFusion (Shen et al. 2024) incorporate Transformer-based RGB+T feature fusion modules for multispectral object detection. In semantic segmentation, EGFNet (Zhou et al. 2022) leverages prior edge maps of RGB+T images to capture detailed contour information. Meanwhile, depth images exhibit valuable insights into spatial structures and 3D layouts. For example, HidaNet (Wu et al. 2023) utilizes a granularity-based attention mechanism to enhance the advantages of RGB+D features for Salient Object Detection (SOD). However, current cross-modal visual understanding is hindered by the lack of well-aligned RGB+X image pairs for training deep models.
@@ 2.2 Layout-to-Image Generation
Based on Denoising Diffusion Probabilistic Model (DDPM) (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020), diffusion models have undergone advancements in training and sampling techniques (Song, Meng, and Ermon 2021; Nichol and Dhariwal 2021). Recent studies have proposed layout-to-image diffusion models, which provide precise instructions for the object placement. Notably, LAW-Diffusion (Yang et al. 2023) integrates a spatial parser and adaptive guidance, enabling complex scene generation. BoxDiff (Xie et al. 2023) directly incorporates spatial conditions into the training-free denoising process. Additionally, InstanceDiffusion (Wang et al. 2024) focuses on precise instance-level control in image generation.
@@ 2.3 Multi-Modal Generation
Multi-modal generation has become a challenging task and attracted widespread research. MM-Diffusion (Ruan et al. 2023) integrates featuring two interconnected denoising autoencoders and a sequential multi-modal U-Net architecture for joint audio-video generation. Moreover, MM-Interleaved (Tian et al. 2024) dynamically extracts information from multi-scale features for image-text generation. In addition, ControlVAR (Li et al. 2024), MT-Diffusion (Chen et al. 2024), and 4M (Mizrahi et al. 2024) adopt large-scale natural datasets or private datasets for end-to-end training on conditional generation. Despite these advancements, there is currently no generative model specifically focused on layout-guided cross-modal “RGB+X” image generation.
3 Cross-Modal Dataset Construction
Since DiffX aims to sample cross-modal images under user instruction, it is crucial to construct datasets with accurate layout control and text description for model training. However, existing cross-modal datasets usually lack detailed text captions. In this section, we introduce the process of constructing the image captions, as illustrated in Figure 2 (a).
Caption Generation: To obtain high-quality image captions, we employ the advanced LMM, namely LLaVA-v1.5-7b model (Liu et al. 2024), to extract detailed descriptions of the images by the given prompt, including weather conditions, environmental settings, surrounding elements, transportation conditions, and identifiable objects.
Caption Correction: Despite the effectiveness of LLaVA in capturing visual content, it can sometimes generate inaccurate captions, especially in complex scenes. Therefore, in cases where the generated captions exhibit minimal errors, we conduct manual corrections to ensure the accuracy. However, instances of significant errors or ambiguous descriptions necessitate further corrections. Under such conditions, we utilize the GPT-4o model (Achiam et al. 2023) to rectify inaccurate information and provide additional descriptions.
By combining the capability of the LLaVA model with the manual correction and assistance of GPT-4o, the captioning process ensures fidelity and accuracy in describing the visual context in the cross-modal datasets. Examples of the data and prepared captions are shown in Figure 2 (b).
@@ 4 DiffX for Cross-Modal Generation
The overall workflow of the DiffX model is shown in Figure 3. The noisy latent $\boldsymbol{z}_{T}$ is randomly initialized, which is adopted for generating $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ by our DiffX-UNet. The denoising process involves embedding the text captions and layout conditions for user instruction.
In the following subsections, we first introduce the preliminaries on Latent Diffusion Model (LDM). Then, we describe the workflow and components of our DiffX in detail.
@@ 4.1 Preliminaries on Latent Diffusion Model
Recently, LDM and its successor SD (Rombach et al. 2022) have pushed the boundaries of diffusion models by using a low-dimensional latent space. Inside LDM, a prior VAE is employed to extract the latent representation $\boldsymbol{z}$ of the ground-truth image $\boldsymbol{I}$. Subsequently, the basic denoising model $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$, normally implemented as a UNet (Ronneberger, Fischer, and Brox 2015) with residual (He et al. 2016) and self-attention (Vaswani et al. 2017) blocks, is adopted to obtain the noise $\boldsymbol{\epsilon}$ in the noisy latent $\boldsymbol{z}_{t}$ at time step $t$. The training objective can be represented as:
@@
where $\boldsymbol{c}$ denotes the conditional input, typically a text embedding generated by a text encoder such as CLIP (Radford et al. 2021) in text-to-image generation.
Following the training process, the denoising model $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$ gradually generates $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ from a randomly initialized noisy latent $\boldsymbol{z}_{T}$. Finally, the decoder of VAE is employed to generate the target image $\boldsymbol{I}^{\prime}$ based on $\boldsymbol{z}_{0}$.
@@ 4.2 Cross-Modal Modeling Formulation
In our DiffX model, we first obtain the joint-modality distribution by mapping the “RGB+X” images into the shared latent space. Specifically, during the forward process, the noising transition from $t-1$ to $t$ is conditioned on the joint modality $\mathbf{M}=\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$, which can be expressed as:
@@
Meanwhile, the reverse process can be regarded as the converse of the forward distributions, resulting in a joint distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1},\mathbf{M}|\boldsymbol{z}_{t})$ at time step $t$:
@@
where $\boldsymbol{\theta}$ denotes the parameters of the DiffX model.
@@ 4.3 Dual-Path Variational AutoEncoder
As illustrated in Figure 4, the proposed DP-VAE is critical for our cross-modal generative task. In contrast to the conventional VAE used in LDM, our DP-VAE employs the input form of “RGB+X” and output form of “RGB+X”. In detail, the DP-VAE utilizes a single encoder $\mathcal{E}$ to encode the input $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$} (denoted as $\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$ in Eq. (2)) into the modality-shared latent representation $z$. The input images are processed through the separate convolutional layers, followed by element-wise addition before being fed to $\mathcal{E}$. Then, it incorporates two parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ to generate the corresponding output $\boldsymbol{I}_{v}^{\prime},\boldsymbol{I}_{x}^{\prime}$}.
In addition, we have observed that image modality “X” usually exhibits distinct contour edges and prominent object positions. To enhance the reconstruction ability of the DP-VAE, we apply a Laplacian Pyramid (LP) (Burt and Adelson 1987) to extract the high-frequency information from the input cross-modal image pairs. Subsequently, the extracted features are embedded into the multi-scale layers of the encoder through cascaded convolutional layers, enabling a comprehensive enhancement at various frequencies.
During the inference phase of DiffX, the parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ from pre-trained DP-VAE are employed to generate the cross-modal “RGB+X” images based on the denoised latent feature $\boldsymbol{z}_{0}$.
@@
4.4 Joint-Modality Embedder
Before the denoising process, it is necessary to embed the layout and text conditions into grounding features. To establish a comprehensive connection between layout and text conditions, we propose the joint-modality embedder ($\mathcal{J}_{\omega}$):
$\boldsymbol{H}^{*},\boldsymbol{c}^{*}=\mathcal{J}_{\omega}(\text{Layout},\text{Text}).$ (4)
where $\boldsymbol{H}^{*}$ is the text-aware layout feature, and $\boldsymbol{c}^{*}$ is the embedded caption feature. In detail, $\mathcal{J}_{\omega}$ is composed of the layout encoder, Long-CLIP, a gated cross-attention layer, and a Feed Forward (FF) layer. The FF layer is an MLP with middle-dimensional expansion.
@@ 4.4 Joint-Modality Embedder
1) For the box-based layout, boxes $\boldsymbol{b}_{i}$ and labels $\boldsymbol{l}_{i}$ are embedded via Fourier mapping $\mathcal{F}(\cdot)$ (Tancik et al. 2020) and CLIP text encoder $f_{\text{clip}}(\cdot)$, respectively. Then, a Multi-Layer Perception (MLP) with parameters $\phi$ is used to encode them into grounding tokens:
@@
2) For the semantic mask layout represented by $\boldsymbol{m}$, we utilize the the pre-trained ConvNeXtψ model (Liu et al. 2022) to extract the in-depth semantic feature. Subsequently, the position embedding $\boldsymbol{p}$ is added, and an MLP with parameters $\varphi$ is used to generate the layout feature:
@@
where $\mathcal{R}$ denotes the process of reshaping and flattening the semantic feature into tokens.
@@
Text conditions $\boldsymbol{c}$ remain equally significant in cross-modal generation as they provide detailed image descriptions, including various aspects such as weather, environments, and transportation. Specifically, our task necessitates long text captions to describe the small objects and complex backgrounds present in the cross-modal scene, as shown in Figure 2 (b). In our DiffX, we utilize the advanced Long-CLIP as the text encoder, which represents notable improvement over its predecessor, CLIP, due to its ability to support text inputs of up to 248 tokens. In detail, each caption $\boldsymbol{c}$ is embedded into a sequence of text embeddings $\boldsymbol{c}^{*}=f_{\text{L-clip}}(\boldsymbol{c})$.
Gated Cross-Attention.
Inside our $\mathcal{J}_{\omega}$, the gated Cross-Attention (CA) layer is implemented to fuse the layout and text conditional features, which is formulated as follows:
$\text{CA}(\boldsymbol{H},\boldsymbol{c}^{*})=\text{softmax}\left(\frac{W_{q}(\boldsymbol{H})\cdot W_{k}(\boldsymbol{c}^{*})^{T}}{\sqrt{d_{k}}}\right)\cdot W_{v}(\boldsymbol{c}^{*}),$ (7)
where $W_{q}$, $W_{k}$, and $W_{v}$ denote the linear transformations, and $d_{k}$ denotes the dimension of $W_{k}(\boldsymbol{c}^{*})$. Here, the gated cross-attention derives the query from layout condition $\boldsymbol{H}$, while deriving key and value from text condition $\boldsymbol{c}^{*}$.
In detail, a gated CA layer and an FF layer are sequentially implemented in $\mathcal{J}_{\omega}$, which aim to generate the text-aware layout conditional feature $\boldsymbol{H}^{*}$:
$\boldsymbol{H}^{\prime}=\boldsymbol{H}+\lambda\cdot\text{tanh}(\gamma_{1})\cdot\text{CA}(\boldsymbol{H},\boldsymbol{c}^{*}),$ (8)
where the hyperparameter $\lambda$ plays a crucial role in balancing feature quality and controllability. Additionally, $\gamma_{1}$ and $\gamma_{2}$ are two adaptive learnable scalars inside the tanh activation.
@@
4.5 DiffX-UNet
As illustrated in Figure 3, DiffX-UNet serves as the basic model for the denoising process in the modality-shared latent space. It is composed of cascaded residual blocks and Spatial-Transformer blocks. The Spatial-Transformer block is the critical component of the DiffX-UNet, consisting of cascaded attention-based layers and FF layers. To enhance the training efficiency, pre-trained model weights from SD-v1-4 are initialized for the frozen layers.
The joint-modality feature $\boldsymbol{H}^{*}$ embedded by $\mathcal{J}_{\omega}$ is fed into the gated Self-Attention (SA) layer within the Spatial-Transformer blocks, which acts as an adapter module to capture the cross-modal relationships. Similar to the gated CA layer, an FF layer is integrated after each gated SA layer to build the long-range interaction:
$\boldsymbol{T}^{\prime}=\boldsymbol{T}+\mu\cdot\text{tanh}(\delta_{1})\cdot\mathcal{T}(\text{SA}([\boldsymbol{T},f_{\text{linear}}(\boldsymbol{H}^{*})])),$ (10)